# Zihan Tan    Suyuan Huang    Guancheng Wan    Wenke Huang    He Li    Mang Ye

### v1

@@ 1 Introduction
From the spatial perspective, due to edge loss, we hypothesize that nodes in subgraph-FL lose label signals from originally nearby labeled nodes. This degradation hampers the ability of GNNs to learn comprehensive semantic knowledge, resulting in poor global performance and reduced generalizability. Correspondingly, we define this phenomenon as Label Signal Disruption (LSD), which naturally exists in subgraph-FL. For verifying, inspired by graph active learning research (Han et al., 2023), we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL. Specifically, SIS evaluates the influence and significance of labels on graphs. In LABEL:figure:sis&specbia(a), we empirically demonstrate that the SIS initially decreases and then increases as the number of partitions increases. Correspondingly, existing methods fail to address LSD and suffer from poor class knowledge. Based on our analysis, we pose the question: I) How can we address label signal disruption in subgraph-FL?
From the spectral perspective, inconsistencies in signal frequencies across subgraphs of clients induce spectral client drifts in the signal transmission schemes of GNNs, thereby undermining the collaboration. To verify this phenomenon, we examine graph spectra across clients and demonstrate the heterogeneity in LABEL:figure:sis&specbia(b). It reveals inconsistent eigenvalue distribution across clients. As a result, GNNs learn distinct signal propagation schemes of subgraphs and optimize in different spectral directions, leading to spectral client drifts and degraded generalizability. Based on the above analysis, we pose the question: II) How can we alleviate spectral client drifts under spectral heterogeneity?
@@ 4.1 Motivation
Signal propagation over graph structures fundamentally underpins the signal transmission paradigm in GNNs. Therefore, rather than focusing solely on challenges arising from static graph structures in FGL, as prior works have done, it is crucial to consider the dynamics of signal propagation in FGL. Specifically, the spatial domain governs the explicit transmission of signals between connected nodes, while the spectral domain captures signal diffusion across different frequency components. Accordingly, we empirically validate the presence of two major challenges: label signal disruption and spectral client drifts, from both the spatial and spectral perspectives.
@@
Motivation of NLIR. Graph data is fragmented across multiple clients in FGL, which inevitably disrupts these label signals, especially as the number of partitions increases. Label signal disruption severs key pathways for propagating label information, resulting in incomplete and biased local feature representations that ultimately degrade GNNs ability semantically. For validation, we empirically investigate how the Structure Inertia Score (SIS) varies in subgraph-FL and experimentally demonstrate the relationship between the decrease in SIS and the client scale. Specifically, the SIS first decreases and then increases as the number of client partitions increases, highlighting the significant role of LSD in the observed performance decline. We aim to mitigate the impact of label signal disruption in FGL by preserving valuable label knowledge across fragmented data. Correspondingly, Node Label Information Reinforcement (NLIR) is introduced. By selecting nodes with both structural representativeness and rich label information for global repository and injecting it during local training, NLIR reduces the information loss inherent in subgraph partitioning. Specifically, throughout the training process, NLIR assesses the similarity distributions between node features and all representative prototypes, thereby enabling multidimensional localization of node classes. It effectively reconstructs label signal pathways within each client, enhancing local feature modeling by restoring context around labeled nodes.
@@
Motivation of FGMA. We reveal spectral client drifts in the subgraph scenario, where GNNs across different clients capture high and low-frequency information inconsistently due to local adaptations. This inconsistency results from frequency-based signal propagation variations, which amplify biases during collaboration and compromise the generalizability of the global model. Spectral client drifts manifest when models optimized for distinct local frequency characteristics are combined, leading to a global model that is unable to generalize well across diverse frequency profiles. To address spectral client drifts, we propose the Frequency-aware Graph Modeling Alignment. Specifically, FGMA constructs the local graph spectra of the GNN adjacency awareness after calculating node similarity matrix. By projecting features separately onto high- and low-frequency components of the reconstructed spectra and aligh the local and global projections, FGMA promotes the generalizability of the global GNN graph signal propagation paradigm, thereby reducing frequency-based discrepancies duringcollaboration.
@@ 4.2 Node Label Information Reinforcement
First of all, we introduce the Structure-Aware Label Centrality (SALC) metric. The SALC, denoted as $\Lambda^{\text{SALC}}_{u}$, is defined as the combination of the Label Influence Centrality $\Lambda^{l}_{u}$ and the Structural Prominence Score $\Lambda^{s}_{u}$:
@@ 5.1 Experimental Setup
$S^{2}$ (ours) $\mathbf{83.4\pm 0.5}$1.5 $\mathbf{76.0\pm 0.3}$1.7 $\mathbf{88.6\pm 0.1}$1.3 $\mathbf{74.7\pm 0.1}$1.9 $78.4\pm 0.1$0.8 $\mathbf{80.5\pm 0.3}$0.9
@@ 5.2 Experiment Results
Varying Scaling Factors of NLIR and FGMA. For the NLIR method, we test scaling factors set to 100, 50, 10, and 1. For the FGMA method, the settings were 0.01, 0.05, 0.5, and 1. The results, shown in LABEL:fig:_hy, indicate that our method maintains consistent performance across these varying hyperparameter configurations.
Varying Client Scales. We assessed performance with different numbers of client partitions: 5, 10 and 20 clients. We compared $S^{2}$ with other methods, including FedAvg , FedProx , and FGSSL. The outcomes, presented in LABEL:fig:_client, demonstrate that $S^{2}$ consistently delivers reliable results regardless of the number of partitions.
@@
Q4: Does the effectiveness of NLIR help alleviate the impact of SIS decline in subgraph-FL? Through experiments, we observe that as the number of clients increases, the Subgraph Information Similarity (SIS) consistently decreases, indicating greater structural divergence across local graphs. Notably, the performance gains brought by NLIR become more pronounced under these conditions. This trend, illustrated in Fig. 5, highlights the targeted effectiveness of NLIR in mitigating the adverse effects of reduced SIS in subgraph-FL settings.
Figure 5: Performance improvement of global GNN using the NLIR method in relation to the variation of the Structure Inertia Score (SIS) across different numbers of client partitions.

### v2

@@ 1 Introduction
Figure 1: (a) According to existing FGL mainstream research, the range of 5 to 20 is the closest to real-world application scenarios. This range is encountering the toughest label signal disruption challenges. (b) The heat map of Kullback-Leibler divergence of eigenvalue distributions across clients. Inconsistency in graph signal frequency of subgraphs leads to spectral client drifts.
@@
From the spatial perspective, due to edge loss, we hypothesize that nodes in subgraph-FL lose label signals from originally nearby labeled nodes. This degradation hampers the ability of GNNs to learn comprehensive semantic knowledge, resulting in poor global performance and reduced generalizability. Correspondingly, we define this phenomenon as Label Signal Disruption (LSD), which naturally exists in subgraph-FL. For verifying, inspired by graph active learning research (Han et al., 2023), we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL. Specifically, SIS evaluates the influence and significance of labels on graphs. In Fig. 1(a), we empirically demonstrate that the SIS initially decreases and then increases as the number of partitions increases. Correspondingly, existing methods fail to address LSD and suffer from poor class knowledge. Based on our analysis, we pose the question: I) How can we address label signal disruption in subgraph-FL?
From the spectral perspective, inconsistencies in signal frequencies across subgraphs of clients induce spectral client drifts in the signal transmission schemes of GNNs, thereby undermining the collaboration. To verify this phenomenon, we examine graph spectra across clients and demonstrate the heterogeneity in Fig. 1(b). It reveals inconsistent eigenvalue distribution across clients. As a result, GNNs learn distinct signal propagation schemes of subgraphs and optimize in different spectral directions, leading to spectral client drifts and degraded generalizability. Based on the above analysis, we pose the question: II) How can we alleviate spectral client drifts under spectral heterogeneity?
@@ 4.1 Motivation
Signal propagation over graph structures fundamentally shapes the signal transmission paradigm in GNNs. Therefore, rather than focusing solely on challenges arising from static graph structures in FGL, as prior works have done, it is crucial to consider the dynamics of signal propagation. Specifically, the spatial domain governs the explicit transmission of signals between connected nodes, while the spectral domain captures signal passing across different frequency components. Accordingly, we empirically validate the presence of two major challenges: label signal disruption and spectral client drifts, from both the spatial and spectral perspectives.
@@
Motivation of NLIR. Graph data is fragmented across multiple clients in subgraph-FL, which inevitably disrupts these label signals, especially as the client scale increases. Label signal disruption severs key pathways for propagating label information, resulting in incomplete and biased local feature representations that ultimately degrade GNNs ability semantically. For validation, we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL and experimentally demonstrate the relationship between the decrease in SIS and the client scale. Specifically, the SIS first decreases and then increases as the number of client partitions increases, highlighting the significant role of LSD in the observed performance decline. We aim to mitigate the impact of label signal disruption in FGL by preserving valuable label knowledge across fragmented subgraphs. Correspondingly, Node Label Information Reinforcement (NLIR) is introduced. By selecting nodes with both structural representativeness and rich label information for a global repository and injecting it during local training, NLIR reduces the information loss inherent in the subgraph scenarios. Specifically, throughout the training process, NLIR assesses the similarity distributions between node features and all representative prototypes for both local and global, thereby enabling multidimensional localization of node classes. By aligning the two similarity distributions, it effectively reconstructs label signal pathways within each client, enhancing local feature modeling by restoring context around labeled nodes.
@@
Motivation of FGMA. We reveal the challenge of spectral client drifts in the subgraph scenario, where GNNs across different clients capture high and low-frequency information inconsistently due to graph structural heterogeneity. This results from overfitting local signal propagation frequencies, which bring biases during collaboration and compromise the generalizability of the global model. To address spectral client drifts, we propose the Frequency-aware Graph Modeling Alignment. Specifically, FGMA reconstructs the local graph spectra by leveraging the GNN adjacency awareness after calculating the node similarity matrix. By projecting features separately onto high- and low-frequency components of the reconstructed spectra and aligh the local and global projections, FGMA promotes the learning of a generalizable spectral signal propagation paradigm, thereby reducing frequency-based discrepancies during collaboration and enhancing the global generalizability.
@@ 4.2 Node Label Information Reinforcement
First of all, we introduce the Structure-Aware Label Centrality (SALC) metric, denoted as $\Lambda^{\text{SALC}}_{u}$. It is defined as the combination of the Label Influence Centrality $\Lambda^{l}_{u}$ and the Structural Prominence Score $\Lambda^{s}_{u}$:
@@ 5.1 Experimental Setup
$S^{2}$ (ours) $\mathbf{83.4\pm 0.2}$1.5 $\mathbf{76.0\pm 0.3}$1.7 $\mathbf{88.6\pm 0.2}$1.3 $\mathbf{74.8\pm 2.3}$2.0 $\mathbf{79.0\pm 1.0}$1.4 $\mathbf{80.5\pm 0.1}$0.9
@@ 5.2 Experiment Results
Varying Scaling Factors of NLIR and FGMA. For the NLIR method, we test scaling factors set to 100, 50, 10, and 1. For the FGMA method, the settings were 0.01, 0.05, 0.5, and 1. The results, shown in Fig. 4, indicate that our method maintains consistent performance across these varying hyperparameter configurations.
(a) Cora
(b) Citeseer
Figure 4: Analysis of the performance growth between $S^{2}$ and FedAvg under varying scaling factors of NLIR and FGMA
Varying Client Scales. We assessed performance with different numbers of client partitions: 5, 10 and 20 clients. We compared $S^{2}$ with other methods, including FedAvg , FedProx , and FGSSL. The outcomes, presented in Fig. 5, demonstrate that $S^{2}$ consistently delivers reliable results regardless of the number of partitions.
(a) Cora
(b) Citeseer
Figure 5: Analysis of performance under different client numbers.
@@
Through experiments, we observe that as the number of clients increases, the SIS consistently decreases, indicating greater structural divergence across local graphs. Notably, the performance gains brought by NLIR become more pronounced under these conditions. This trend, illustrated in Fig. 6, highlights the targeted effectiveness of NLIR in mitigating the adverse effects of reduced SIS in subgraph-FL settings.
Figure 6: Performance improvement using the NLIR method in relation to the variation of the SIS across different client scales.